In the graph on the right, the demand for syrup has changed because the price of frozen waffles has risen from $2.00 to $2.50 per package. Syrup 3.7 The cross-price elasticity of demand between frozen waffles and syrup is(Use the midpoint formula and enter your response rounded to two decimal places. Be sure to include the minus sign if necessary.) 2.7 2.00-1.89 1.5 : 7,0009,000 00 6000 9,000 12.000 15,000 Quantity (packages of syrup per week)
Answer #1

Frozen waffles and syrup are complementary goods. If price of frozen waffles rises from $2 to $2.50. Then , demand for syrup shifts leftward from D1 to D2 , as a result quantity demanded of syrup decreases from 9000 to 7000 packages.

Cross price elasticity of demand = (% change in quantity demanded of syrup / % change in price of frozen waffles)

By mid-point method:

% change in quantity demanded of syrup = [(7000-9000)/ [(7000+9000)/2] ] 100 =[(-2000)/ (8000)] 100 = -25%

% change in price of frozen waffles = [(2.50-2) / [(2+2.50)/2] ] 100= (0.50 /2.25) 100= 22.22%

Cross price elasticity = (-25%)/ (22.22%)= -1.13